OpenStudy (anonymous):

“numbers that are at least 6 units but no more than 10 units from 2

OpenStudy (anonymous):

\[6 \le\left| x-2 \right| \le 10\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

In general AbsValue(a - b) represents the "distance of a from b"